### v2

# LLM-based NLG Evaluation: Current Status and Challenges

Abstract
Evaluating natural language generation (NLG) is a vital but challenging problem in artificial intelligence. Traditional evaluation metrics mainly capturing content (e.g. n-gram) overlap between system outputs and references are far from satisfactory, and large language models (LLMs) such as ChatGPT have demonstrated great potential in NLG evaluation in recent years. Various automatic evaluation methods based on LLMs have been proposed, including metrics derived from LLMs, prompting LLMs, and fine-tuning LLMs with labeled evaluation data. In this survey, we first give a taxonomy of LLM-based NLG evaluation methods, and discuss their pros and cons, respectively. We also discuss human-LLM collaboration for NLG evaluation. Lastly, we discuss several open problems in this area and point out future research directions.
**footnotetext: Equal contribution.
1 Introduction
Evaluating natural language generation (NLG) is a key but challenging issue. Traditional evaluation metrics like BLEU Papineni et al. (2002) and ROUGE Lin (2004) rely on the n-gram overlap between model outputs and references to measure its quality. They have been criticized for low correlation with human judgments Sulem et al. (2018), as surface-level matching cannot reliably evaluate text. After the rise of deep learning, model-based evaluation metrics like BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021) have been continuously proposed and gradually adopted to evaluate the overall quality or various specific aspects of generated outputs (e.g., fluency, coherence, coverage, faithfulness, etc.). Although better than traditional metrics, their performance is still not satisfactory, and their application scope is very limited. For example, BERTScore is reference-based and cannot be used without a reference. With the emergence of large language models (LLMs) like InstructGPT Ouyang et al. (2022), they have achieved unprecedented effectiveness in following instructions, understanding content, and generating text. This inspired researchers to use LLMs for NLG evaluation. Although this is a research direction that only emerged in 2023, the past year has seen an enormous amount of research work. It is no exaggeration to say that NLG evaluation has been revolutionized by LLMs. This article will review the existing literature and provide suggestions for future research in this field.
Figure 1: Schematic representation of the four categories of LLM-based NLG evaluation.
This article mainly focuses on research that uses language models with over one billion parameters for NLG evaluation, with necessary references to earlier model-based evaluation metrics like BERTScore. To maintain focus, other types of generation like code generation and tasks involving images are not included in the scope of this article. As shown in Figure 1, according to how we utilize LLMs for NLG evaluation, we categorize the research work into four types:
  • LLM-derived Metrics2): developing/deriving evaluation metrics from embeddings or generation probabilities of LLMs.
  • Prompting LLMs3): directly inquiring existing LLMs via designed prompts which involve different elements for evaluation.
  • Fine-tuning LLMs4): using labeled evaluation data to fine-tune existing LLMs and improving their NLG evaluation capabilities.
  • Human-LLM Collaborative Evaluation5): leveraging distinctive strengths of both human evaluators and LLMs to achieve robust and nuanced evaluations in challenging domains through human-LLM collaboration.
We will review each type of evaluation methods and discuss the pros and cons respectively. Lastly, we will discuss future directions in this area (§6).
2 LLM-derived Metrics
2.1 Overview
Early model-based NLG evaluation methods and metrics, such as BERTScore and BARTScore, were primarily motivated by the capability of traditional pre-trained language models to generate high-quality texts. Recently, the advent of LLMs has prompted some research to adapt these ideas to stronger LLMs. Their inherent, powerful linguistic abilities are derived for direct NLG evaluation, which is expected to result in better performance. Such works can be categorized into two main types: embedding-based metrics and probability-based metrics.
2.2 Embedding-based Metrics
The embedding-based methods, like BERTScore, generally utilize representations of language models and thus compute the semantic similarity between the reference and the target text to evaluate, with different possible ways of implementation. Recent work ES et al. (2023) similarly obtains embeddings for the target text using the text-embedding-ada-002 model through the OpenAI API. The higher the corresponding similarity, the closer the target text aligns with the relevant requirements, indicating a higher quality.
2.3 Probability-based Metrics
To better utilize the knowledge inherent in language models, probability-based methods like BARTScore formulate text generation evaluation as conditional probability comparison, positing that the better the quality of the target text, the higher the likelihood that models should be able to generate it. Recently, GPTScore Fu et al. (2023a) has established tailored evaluation templates for each aspect to effectively guide multiple LLMs for NLG evaluation, including GPT3 Brown et al. (2020), OPT Zhang et al. (2022), and FLAN Chung et al. (2022). The generation probability is calculated under the condition of source input designed with customized prompts and evaluation aspects, which endows the GPTScore with better flexibility in evaluation. And similar methods have also been applied to the hallucination detection of the LLM-generated text Varshney et al. (2023) with three different attempts for calculating the probability score.
On the other hand, some works leverage the variation in probabilities under changed conditions as the evaluation metric. FFLM Jia et al. (2023) proposes to evaluate the faithfulness of the target text by calculating a combination of probability changes based on the intuition that the generation probability of a given text segment increases when more consistent information is provided, and vice versa. Similarly, DELTASCORE Xie et al. (2023) measures the quality of different story aspects according to the likelihood difference between pre- and post-perturbation states with LLMs including GPT-3.5 (text-davinci-003) that provide logits. They believe that the sensitivity to specific perturbations indicates the quality of related aspects, and their experiments demonstrate the effectiveness of their approach.
2.4 Pros and Cons
Traditional NLG evaluation approaches always fall short due to their surface-form similarity when the target text and reference convey the same meaning but use different expressions. In contrast, LLM-derived metrics offer a remedy for the limitation and demonstrate stronger correlations with human judgments benefiting from the evolving modeling techniques. However, the flaws within LLMs can lead to some issues, as introduced in the following:
Robustness. Some research has investigated the robustness of LLM-derived metrics and found that they lack robustness in different attack scenarios. Specifically, He et al. (2023b) develops a set of stress tests to assess the robustness of various model-based metrics on some common NLG tasks. They show a catalogue of the blind spots and potential errors identified that are not detected by different metrics.
Efficiency. Compared to traditional metrics, LLM-derived evaluation methods are more time-consuming and require more computational resources, especially when adopting LLMs with quite large parameter scales. To address this, Eddine et al. (2022) proposes an approach to learning a lightweight version of LLM-derived metrics, and some fast LLM inference and serving tools like vLLM Kwon et al. (2023) have been launched. However, closed-source LLMs often do not make their parameters, representations, or logits public and available, thus making it impossible to apply LLM-derived methods to them.
Fairness. Sun et al. (2022) assesses the social bias across various metrics for NLG evaluation on six sensitive attributes: race, gender, religion, physical appearance, age, and socioeconomic status. Their findings reveal that model-based metrics carry noticeably more social bias than traditional metrics. Relevant biases can be categorized into two types: intrinsic bias encoded within pre-trained language models and extrinsic bias injected during the computation of similarity. Therefore, current LLM-derived methods may have similar issues.
3 Prompting LLMs
3.1 Overview
LLMs have demonstrated unprecedented instruction understanding and text generation abilities, which broadens researchers’ imagination of automatic evaluation for NLG. For a long time, human evaluation has been viewed as the gold standard for NLG evaluation. Recently, some studies claim that LLMs are on par with crowdsourcing annotators in several tasks Törnberg (2023); Gilardi et al. (2023); Ostyakova et al. (2023); Cegin et al. (2023). So, can LLMs simulate or even be an alternative to humans in the human evaluation of NLG? Or, do the practices in human evaluation or other evaluative tasks (e.g. competition race, paper review, etc.) inspire us to better NLG evaluation with LLMs? A large body of research and attempts based on prompting LLMs are guided by these ideas. In these studies, instructions and the text to be evaluated are completely expressed in the prompt given to LLMs, and the evaluation results are generated by them.
Human evaluation typically includes the following elements:
  • Evaluation Methods: The way the preferences of annotators are obtained, such as scoring and comparison.
  • Task Instructions: How the annotators should read or manipulate different parts to complete the annotation.
  • Input Content: The target text to be evaluated and other required content. Other required content including source documents, references, and external knowledge is provided as needed.
  • Evaluation Criteria: The general definition of how good or bad the text to be evaluated is in a particular aspect of quality, e.g. fluency, faithfulness.
  • Role and interaction: The roles annotators play in the evaluation and the interactions between them.
The focus of the existing research can always be mapped analogously to one or more of these elements, and we organize them according to the elements they address. An example of prompting LLMs is shown in Figure 2.
Figure 2: An example of prompting LLMs to evaluate the aspect of consistency of the summary. There are task instructions, evaluation criteria, input content, and evaluation methods in the prompt, as well as the evaluation results, including the rating and explanation generated by LLMs.
3.2 Evaluation Methods
Scoring. Scoring is the most commonly used evaluation method in human evaluation for NLG, which is naturally applied to NLG evaluation by prompting LLMs. Chiang and Lee (2023a) has conducted relevant studies early, using a Likert scale from 1 to 5 to evaluate story generation and adversarial attacks with InstructGPT Ouyang et al. (2022) and ChatGPT  [^1] https://openai.com/blog/chatgpt/, showing that the evaluation results of LLMs are consistent with expert human evaluators. Kocmi and Federmann (2023b) discover GPT-3.5 and GPT-4 achieve the state-of-the-art accuracy of evaluating translation quality compared to human labels, outperforming all the results from the metric shard task of WMT22 Freitag et al. (2022). Wang et al. (2023a) experiment on five datasets across summarization, story generation, and data-to-text, and ChatGPT evaluators with a rating scale from 1 to 5 or 1 to 100 have the state-of-the-art or comparative correlations with human judgments in most settings, compared with prior metrics. Similar conclusions are also observed in open-domain dialogue response generation Lin and Chen (2023). Besides English, Mendonça et al. (2023) show that ChatGPT with simple rating prompts is a strong evaluator for multilingual dialogue evaluation, surpassing prior metrics based on encoders.
Comparison. Different from absolute scoring, comparison refers to choosing the better of the two. Luo et al. (2023); Gao et al. (2023) use ChatGPT to compare the factual consistency of two summaries. AuPEL Wang et al. (2023d) evaluate personalized text generation from three aspects in the form of comparison with the PaLM 2 family Anil et al. (2023). According to Liusie et al. (2023), pairwise comparison is better than scoring when medium-sized LLMs (e.g. FlanT5 Chung et al. (2022) and Llama2 Touvron et al. (2023)) are adopted as evaluators.
Ranking. Ranking can be viewed as an extended form of comparison. In comparison, only two examples are involved at a time, whereas in ranking, the order of more than two examples needs to be decided at once. Ji et al. (2023) use ChatGPT to rank five model-generated responses across several use cases at once, indicating the ranking preferences of ChatGPT align with those of humans to some degree. Similarly, GPTRank is a method to rank summaries in a list-wise manner Liu et al. (2023c). Moreover, Liu et al. (2023b) compare different evaluation methods in LLM-based summarization including scoring, comparison, and ranking, showing that the optimal evaluation method for each backbone LLM may vary.
Boolean QA. Boolean QA requires LLMs to answer ”Yes” or ”No” to a question. It is adopted more in scenarios where human annotations are binary, such as grammaticality Hu et al. (2023), faithfulness of summaries and statements Luo et al. (2023); Gao et al. (2023); ES et al. (2023); Hu et al. (2023), factuality of generated text Fu et al. (2023b); Guan et al. (2023); Manakul et al. (2023), and answerability of generated questions Wang et al. (2023f).
Error Analysis. Error Analysis refers to the evaluation of a text by looking for errors that occur in the text according to a set of predefined error categories. Multidimensional Quality Metrics (MQM) Jain et al. (2023) is an error analysis framework prevalent in machine translation evaluation. According to MQM, Lu et al. (2023); Kocmi and Federmann (2023a) use ChatGPT or GPT-4 to automatically detect translation quality error spans. BOOOOKSCORE Chang et al. (2023), an LLM-based evaluation metric, assesses the coherence of book summaries by identifying eight types of errors.
3.3 Task Instructions
In human evaluation, this part usually comes in the form of a task description or evaluation steps. They can also exist at the same time. The task description states the annotation in a more general way, and the evaluation steps, which can be considered as Chain-of-Thought, explicitly describe what to do at each step. In the context of prompting LLMs for NLG evaluation, the demonstrations used in few-shot prompting are also included.
Form and requirements. Several studies from an Eval4NLP 2023 shared task Leiter et al. (2023) have explored task instructions in various settings. Kim et al. (2023a) conduct experiments on different templates and lengths of task descriptions and evaluation steps. Kotonya et al. (2023) generate task instructions with LLMs or improve existing task instructions with LLMs. Additionally, He et al. (2023a) evaluate generative reasoning using LLMs by asking them first to generate their own answers, and then conduct a quantitative analysis of the text to be evaluated.
Analysis and explanations. LLMs can include analysis or explanation in their evaluations, which is a key point that distinguishes them from previous automatic evaluation metrics. Early explorations into prompting LLMs for NLG evaluation mostly do not examine the impact of whether LLMs are required to analyze and explain on evaluation results. However, Chiang and Lee (2023b) explore different types of evaluation instructions in summarization evaluation and dialogue evaluation, finding that explicitly asking large models to provide analysis or explanation achieve higher correlation with human judgments. Besides, the quality of the analysis and explanation generated by LLMs itself requires additional manual evaluation Leiter et al. (2023). Naismith et al. (2023) compare the explanations written by humans and generated by GPT-4 and conduct a simple corpus analysis on the generated explanations.
In-context examples. Similarly to other fields, sometimes demonstrations are needed when prompting LLMs for NLG evaluation. Specifically, Jain et al. (2023) use only in-context examples as task instructions, relying on LLMs to evaluate the quality of summaries. In scenarios where task descriptions or evaluation steps are included, Kotonya et al. (2023) compare the performance of LLMs as evaluators in both zero-shot and one-shot settings, finding that one-shot prompting does not bring improvements. Moreover, Hasanbeig et al. (2023) improve the performance of LLM evaluators by updating the in-context examples iteratively.
3.4 Input Content
The types of input content mainly depend on the evaluation criteria and are relatively fixed. For most task-specific evaluation criteria, such as the faithfulness of a summary Luo et al. (2023); Gao et al. (2023), the source document is needed in addition to the target text to be evaluated. For task-independent criteria, such as fluency Hu et al. (2023); Chiang and Lee (2023b), only the text to be evaluated needs to be provided, though many works also provide the source document Wang et al. (2023a); Liusie et al. (2023). Other types of input content can be provided as required by the specific task. Kocmi and Federmann (2023b) use two different settings when evaluating machine translation: providing references and not providing references. Guan et al. (2023) provide relevant facts and context when evaluating whether a text conforms to the facts. Exceptionally, Shu et al. (2023) add the output of other automatic evaluation metrics to the input of the LLM.
3.5 Evaluation Criteria
The evaluation targeting specific aspects is used in numerous studies of human evaluation for NLG, such as text summarization, story generation, dialogue, and text simplification. Evaluation criteria, i.e., the definitions of aspects are key in this context. Most evaluation criteria in LLM-based evaluation are directly derived from human evaluation. However, a few studies have attempted to let LLMs generate or improve evaluation criteria. Liu et al. (2023e) use a few human-rated examples as seeds to let LLMs draft some candidate evaluation criteria, and then further filter them based on the performance of LLMs using these criteria on a validation set, to obtain the final evaluation criteria. Kim et al. (2023c) designed an LLM-based interactive evaluation system, which involves using LLMs to review the evaluation criteria provided by users, including eliminating ambiguities in criteria, merging criteria with overlapping meanings, and decomposing overly broad criteria. Additionally, Ye et al. (2023a) propose a hierarchical aspect classification system with 12 subcategories, demonstrating that under the proposed fine-grained aspect definitions, human evaluation and LLM-based evaluation are highly correlated. Besides, the chain-of-aspects approach improves LLMs’ ability to evaluate on a specific aspect by having LLMs score on some related aspects before generating the final score Gong and Mao (2023).
3.6 Role and Interaction
We include in this section the evaluation strategies that either use the same LLMs in different ways or involve different LLMs. The former can be further divided into chain-style and network-style interactions.
Chain-style interaction. Inspired by human evaluators, Yuan et al. (2024) have LLMs score a batch of examples to be evaluated each time. Specifically, the evaluation process is divided into three stages: analysis, ranking, and scoring. Similar to QA-based evaluation metrics Durmus et al. (2020), Fu et al. (2023b) assess the faithfulness of summaries in two stages: treating LLMs as question generators to generate a question from the summary; then having LLMs answer the question using the source document. Differently, when Hu et al. (2023) use GPT-4 to evaluate the faithfulness of summaries, it first asks GPT-4 to extract event units from the summary, then verifies whether these event units meet the requirements, and finally judges whether the event units are faithful to the source document.
Network-style interaction. Unlike chain-style interactions, network-style interactions involve the dispersion and aggregation of information. In network-style interactions, LLMs on the same layer play similar roles. ChatEval Chan et al. (2023) is a framework for evaluating content through debates among multiple LLMs, with three communication strategies designed among the three types of LLMs: One-By-One, Simultaneous-Talk, and Simultaneous-Talk-with-Summarizer. Zhang et al. (2023b) find that under certain conditions, widening and deepening the network of LLMs can better align its evaluation with human judgments. Saha et al. (2023) propose a branch-solve-merge strategy, assigning LLMs the roles of decomposing problems, solving them, and aggregating answers, thereby improving the accuracy and reliability of evaluations. Wu et al. (2023) assume that different people such as politicians and the general public have different concerns about the quality of news summaries, use LLMs to play different roles in evaluation accordingly, and aggregate the results finally.
Different LLMs. Different from having the same LLM play different roles, some research has used different LLMs (such as GPT-4 and Claude) in their studies. In pairwise comparisons, previous work mostly used a single LLM as the evaluator, which may not be fair. In light of this, Bai et al. (2023) design a decentralized Peer-examination method, using different LLMs as evaluators and then aggregating the results. Further, Li et al. (2023c) let different LLMs serve as evaluators in pairwise comparisons and then have them go through a round of discussion to reach the final result. Additionally, Cohen et al. (2023) evaluate the factuality of texts through the interaction of two LLMs, where the LLM that generated the text acts as the examinee and the other LLM as the examiner.
3.7 Pros and Cons
The benefits of prompting LLMs for NLG evaluation are exciting. First, for the first time, people can express evaluation criteria and evaluation methods in natural language within the prompts given to LLMs, providing great flexibility. Where previously people needed to design specific evaluation metrics for different NLG tasks or even different aspects of a single task, now they only need to modify the prompts for LLMs. Secondly, surprisingly, LLMs have the ability to generate explanations while assessing texts, making this approach somewhat interpretable. Furthermore, in many NLG task, prompting LLMs for evaluation has achieved state-of-the-art correlations with human judgments.
However, as many studies have pointed out, this type of approach still has many limitations. Wang et al. (2023b) note that when using ChatGPT and GPT-4 for pairwise comparisons, the order of the two texts can affect the evaluation results, which is known as position bias. To alleviate this issue, Li et al. (2023d) propose a strategy of splitting, aligning, and then merging the two texts to be evaluated into the prompt. Also, LLM evaluators tend to favor longer, more verbose responses Zheng et al. (2023) and responses generated by themselves Liu et al. (2023a). Wu and Aji (2023) show that compared to answers that are too short or grammatically incorrect, answers with factual errors are considered better by LLMs. Liu et al. (2023d) demonstrate through adversarial meta-evaluation that LLMs without references are not suitable for evaluating dialogue responses in closed-ended scenarios: they tend to score highly on responses that conflict with the facts in the dialogue history. Zhang et al. (2023a) also present the robustness issues of LLMs in dialogue evaluation through adversarial perturbations. Shen et al. (2023) indicate that LLM evaluators have a lower correlation with human assessments when scoring high-quality summaries. In addition, Hada et al. (2023) state that LLM-based evaluators have a bias towards high scores, especially in non-Latin languages like Chinese and Japanese. Beyond these shortcomings of performance, both ChatGPT and GPT-4 are proprietary models, and their opacity could lead to irreproducible evaluation results.
4 Fine-tuning LLMs
4.1 Overview
As mentioned above, despite the exciting performance of prompting LLMs like ChatGPT and GPT-4 for NLG evaluation, several shortcomings in practice are inevitable, such as high costs, possibly irreproducible results, and potential biases in LLMs. In response, recent research has shifted towards fine-tuning smaller, open-source LLMs specifically for evaluation purposes, aiming to achieve performance close to GPT-4 in NLG evaluation. Representative works of this type include PandaLM Wang et al. (2023e), Prometheus Kim et al. (2023b), Shepherd Wang et al. (2023c), TIGERScore Jiang et al. (2023), INSTRUCTSCORE Xu et al. (2023), Auto-J Li et al. (2023a), CritiqueLLM Ke et al. (2023) and JudgeLM Zhu et al. (2023). Their main ideas are similar, involving the elaborate construction of high-quality evaluation data, followed by accordingly fine-tuning open-source base LLMs. Nevertheless, there are certain discrepancies in the designs across different works, such as the usage of references and evaluation criteria. We have summarized the key different components of these methods in Table 1 for comparison, which will be elaborated on next.
Method Data Construction Evaluation Method Base LLM Reference Required
Instruction Source Annotator Scale Result Mode Details Specific Criteria
PandaLM Alpaca 52K GPT-3.5 300K Comparison Reason & Reference Unified LLaMA 7B No
Prometheus GPT-4 Construction GPT-4 100K Scoring Reason Explicit LLaMA-2-Chat 7B & 13B Yes
Shepherd Community Critique Data & 9 NLP Tasks Data Human 1317 Overall Judgement Error Identifying & Refinement Unified LLaMA 7B No
TIGERScore 23 Distinctive Text Generation Datasets GPT-4 48K MQM Error Analysis Implicit LLaMA-2 7B & 13B No
INSTRUCTSCORE GPT-4 Construction GPT-4 40K MQM Error Analysis Implicit LLaMA 7B Yes
AUTO-J Real-world User Queries from Preference Datasets GPT-4 4396 Scoring & Comparison Reason Implicit LLaMA-2-Chat 13B No
CritiqueLLM AlignBench & ChatGPT Augmentation GPT-4 9332 Scoring Reason Unified ChatGLM-2 6B, 12B & 66B Flexible
JudgeLM GPT4All-LAION, ShareGPT Alpaca-GPT4 & Dolly-15K GPT-4 100K Scoring & Comparison Reason Unified Vicuna 7B, 13B & 33B Flexible
Table 1: Comparison of the different key components among the representative methods of fine-tuning LLMs.
4.2 Data Construction
Diverse data with high-quality annotations is crucial for the fine-tuning of evaluation models, which mainly involves task scenarios, inputs, target texts to evaluate, and evaluation results. Early NLG evaluation research primarily focused on conventional NLG tasks, such as summarization and dialogue generation. Thus, the task scenarios, inputs, and target texts refer to the corresponding NLP task, source inputs of the task, and outputs generated by specialized systems based on task requirements, respectively. And mainstream datasets for these tasks predominantly employ human annotators to provide evaluation results, which are often considered reliable.
With the recent rise of LLMs, the spectrum of NLG tasks has been broadened to scenarios of instruction and response that are more aligned with human needs. Traditional tasks like summarization with corresponding source inputs can be viewed as kinds of instructions and requirements. Meanwhile, responses generated by various general LLMs generally serve as the target texts now and require more flexible evaluation so that the performance of different LLMs can be compared, promoting further developments. Therefore, to keep pace with the current advancement of modeling techniques, most evaluation methods have adopted the similar instruction-response scenario.
The primary differences in these works actually lie in the construction of instructions, with the purpose of improving either diversity or reliability for the better generalization ability of the fine-tuned model. PandaLM and JudgeLM entirely sample from common instruction datasets, such as Alpaca 52K, while CritiqueLLM adopts small-scale sampling followed by ChatGPT augmentation. In contrast, Prometheus and INSTRUCTSCORE rely on GPT-4 to generate all the instructions based on seed data, whereas Auto-J and Shepherd use real-world data. Moreover, since large-scale human annotation is impractical, most works utilize GPT-4 as the powerful annotator, except for PandaLM and Shepherd, which use GPT-3.5 and human annotation on small-scale community data, respectively. During the construction, they basically all design detailed prompts or guidance and apply heuristic filtering strategies and post-processing methods to mitigate noise. Overall, despite the possible higher quality of human annotation, the corresponding drawback is the difficulty in constructing large-scale datasets, which in turn may hinder adequate model training, while using LLMs for construction is the opposite situation.
4.3 Evaluation Method
As with prompting LLMs, the evaluation methods adopted in these works are highly diversified, involving different evaluation criteria, result modes, and usages of the reference. Given that current instruction-response scenarios encompass different types of tasks, it is unsuitable to specify unified evaluation criteria as in traditional NLG tasks. However, some works still do it this way, while some other methods let LLM annotators adaptively and implicitly reflect the required criteria in their evaluations, like PandaLM, TIGERScore, and AUTO-J. In particular, AUTO-J has meticulously crafted 332 evaluation criteria, matched to different tasks. Furthermore, Prometheus explicitly incorporates evaluation criteria into the inputs of the model, expecting flexible evaluation based on various customized criteria.
More details about the evaluation methods are shown in Table 1. All the works require models to provide detailed information, such as reasons for their evaluation results. And the MQM mode can achieve more informative error analysis, offering stronger interpretability. Moreover, some works do not necessarily require references and then have greater value in practice. And a more optimal method is to concurrently support both reference-based and reference-free evaluations as JudgeLM and CritiqueLLM.
4.4 Fine-tuning Implementation
The fine-tuning process is uniformly implemented by different works on their selected open-source LLMs, like LLaMA, and respective constructed data, with some targeted settings. Specifically, Prometheus maintains balanced data distributions during fine-tuning, including the length and label. JudgeLM eliminates potential biases by randomly swapping sample pairs to be compared and randomly removing references. INSTRUCTSCORE utilizes GPT-4 to provide error annotations for the intermediate outputs of the fine-tuned model for further supervised reinforcement. And based on some preliminary experiments and manual analysis, TIGERScore determines appropriate ratios of different types of data during fine-tuning, which are claimed to be crucial by them. Moreover, CritiqueLLM implements separately, with and without references, and explores the effects of data and model scale. Compared to the vanilla fine-tuning setting, these methods have improved the efficiency of model training and the robustness of evaluations.
4.5 Pros and Cons
Those shortcomings of prompting LLM for NLG evaluation can be significantly alleviated due to the customized implementation of data construction and model fine-tuning here. For instance, most fine-tuned models range between 7B and 13B in the scale of parameters, facilitating low-cost inference use and good reproducibility, with performance close to GPT4 in NLG evaluation. And specific measures can be adopted to prevent related biases found in GPT4 during different stages. Furthermore, this type of approach allows for continuous iteration and improvement of the model to address potential deficiencies or emerging issues discovered in future applications.
However, some biases associated with GPT4 may still persist, as the data construction of most methods employs GPT4 for critical evaluation annotation. On the other hand, the base open-source LLMs selected by existing works are primarily the series of LLaMA. With the rapid updates and improvements of open-source large models recently, it adheres to the intuition that employing a more powerful base LLM should lead to better evaluation performance. However, this means repetitive fine-tuning processes and computational expenses from scratch since directly migrating existing fine-tuned models to the new base LLM is difficult.
Additionally, although many existing methods aspire for more flexible and comprehensive evaluation through fine-tuning, demanding excessive evaluation settings may ultimately lead to poor performance or failure in model training, as found by AUTO-J and CritiqueLLM on criteria and references, respectively. However, there are some disagreements here since Prometheus and JudgeLM show different results. Moreover, considering the different evaluation settings in existing works, it is challenging to conduct a horizontal comparison among them. These issues require further exploration in future research.
5 Human-LLM Collaborative Evaluation
5.1 Overview
While LLMs demonstrate robust evaluation capabilities, there exists a need for further enhancement in terms of their reliability, particularly in establishing a stronger correlation with human evaluation outcomes. Although human evaluation is the gold-standard evaluation approach in NLG, it is recognized for its associated high costs and susceptibility to subjective biases van der Lee et al. (2021); Deriu et al. (2021); Li et al. (2023b). The robust and comprehensive capabilities exhibited by LLMs underscore considerable potential for the development of collaborative evaluation methodologies that integrate both human and LLMs. In recent investigations, researchers have initiated the exploration of collaborative evaluation paradigms, which include traditional NLG evaluation methods such as scoring and explaining Zhang et al. (2021); Li et al. (2023b), broader evaluation methods such as testing and debugging Ribeiro and Lundberg (2022), and auditing NLG models to ensure fairness Rastogi et al. (2023). Furthermore, scholars Saunders et al. (2022) are actively engaging in efforts to address the intricate challenge of scalable oversight Amodei et al. (2016) through the collaboration of humans and LLMs. The objective is to devise strategies for effectively evaluating models in tasks that pose inherent difficulties for human assessors. This collaborative approach seeks to leverage the distinctive strengths of both human evaluators and sophisticated language models to achieve robust and nuanced evaluations in challenging domains.
5.2 Scoring and Explaining
Automated evaluation frequently exhibits a limited correlation with human judgments, while human evaluation, though reliable, is labor-intensive. Zhang et al. (2021) present a human-machine collaborative framework (HMCEval) which conceptualizes dialogue evaluation as a sample assignment problem to ensure the reliability of evaluation outcomes while minimizing human effort and achieves 99% accuracy with half human effort. Recently, LLMs have emerged as a cost-effective alternative to human evaluations. However, both humans and LLMs have limitations, including inherent subjectivity and unreliable judgments, especially in open-ended tasks with diverse requirements.
To address challenges associated with inconsistent evaluation criteria in open-ended tasks and explore synergy between humans and LLM-based evaluators, Li et al. (2023b) proposes a Collaborative Evaluation pipeline (COEVAL), which involves designing a checklist of task-specific criteria and conducting detailed evaluations where LLMs generate initial ideation and humans engage in scrutiny. Depending solely on score predictions is insufficient for ensuring reliable evaluation and error detection, particularly when specific criteria demand nuanced analysis beyond straightforward scoring. Building upon recent developments in explainable NLP (Yin and Neubig (2022); Jung et al. (2022); Ribeiro and Lundberg (2022); Ye et al. (2023b)), COEVAL is assigned the additional task of generating explanations to elucidate evaluation outcomes to facilitate a trustworthy collaborative evaluation process. Results indicate COEVAL effectively evaluates lengthy texts by utilizing LLMs, saving significant time and reducing human evaluation outliers. Despite the involvement of LLMs, human scrutiny remains essential, contributing to the revision of around 20% of LLM evaluation scores for enhanced reliability.
5.3 Broader Evaluation Tasks
The broader evaluation of NLG models involves testing and debugging the models. Current methods often rely on highly variable human creativity and extensive manual effort or are limited to addressing a very specific class of bugs. Ribeiro and Lundberg (2022) introduce AdaTest, a process that uses LLMs in collaboration with human feedback to automatically generate unit tests that highlight bugs in a target model, which proves to make users 5-10 times more effective at identifying bugs and assists users in effectively fixing bugs without introducing new ones. Moreover, LLMs have shown biases and irresponsible behavior, necessitating thorough auditing before deployment Blodgett et al. (2020); Jones and Steinhardt (2022). AdaTest++ Rastogi et al. (2023) draw on insights from literature on human-AI collaboration and sensemaking, and engage with research experts in safe and fair AI, which emphasizes the importance of sensemaking and effective communication between humans and AI to capitalize on their complementary strengths in collaborative auditing. AdaTest++ successfully leverages human strengths, such as schematization and hypothesis testing. Moreover, users identified a range of failure modes across 26 different topics in issues that were revealed in formal audits and those that were previously under-reported. Additionally, ensuring trustworthiness in LLMs for challenging tasks Chen et al. (2021); Nakano et al. (2021); Li et al. (2022); Menick et al. (2022) poses a crucial challenge. Scalable oversight Amodei et al. (2016) aims to effectively evaluate models on tasks challenging for humans and suggests the use of AI for assistance. Saunders et al. (2022) explored providing critiques of model outputs as a form of assistance, demonstrating that model-generated critiques assist humans in identifying overlooked flaws.
5.4 Pros and Cons
The advantages of human-AI collaborative evaluation lie in achieving a balance between efficiency and cost, as demonstrated by COEVAL Li et al. (2023b) achieving this equilibrium. Additionally, there are complementary strengths between humans and AI. For instance, AdaTest++ Rastogi et al. (2023) empowers users to consistently utilize their strengths throughout the auditing process, benefiting significantly from LLM. Users who generate the most topics heavily rely on LLM suggestions while employing their contextual reasoning and semantic understanding to update their mental models vigilantly and identify model failures.
However, there are drawbacks. The evaluation results of LLMs may be sensitive to the formats used to query the model and might require additional support for prompt writing Li et al. (2023b); Rastogi et al. (2023). Furthermore, the current capability to assess confidence levels is not strong enough, making it challenging to determine when to trust the LLM. Furthermore, certain level of human supervision is still necessary, making it less convenient and cost-effective compared to fully automated evaluation.
6 Conclusions and Future Trends
Through the above review of studies on NLG evaluation based on LLMs, we find that these four categories of approaches have their respective strengths and weaknesses, and most of the existing work is concentrated on prompting LLMs. In view of this, we offer some suggestions for future directions in this field.
Unified benchmarks for LLM-based NLG evaluation approaches. As mentioned above, each of the studies that fine-tuned LLMs to construct specialized evaluation models uses different settings and data during testing, making them incomparable. In the research on prompting LLMs for NLG evaluation, there are some publicly available human judgments on the same NLG task, such as SummEval for summarization. However, the existing human judgments have many problems. Firstly, most of the existing data only involve one type of NLG task and a single human evaluation method (e.g., scoring), making it difficult to evaluate LLMs’ performance on different tasks, as well as using different evaluation methods on the same task. Secondly, many of the texts in these human judgments are generated by outdated models (such as Pointer Network) and do not include texts generated by more advanced LLMs. Lastly, many human evaluation datasets are too small in scale. There is an urgent need for large-scale, high-quality human evaluation data covering various NLG tasks and evaluation methods as a benchmark.
NLG evaluation for low-resource languages and new task scenarios. Almost all existing research focuses on English data. However, it is doubtful whether LLMs have similar levels of NLG evaluation capability for texts in other languages, especially low-resource languages. As Zhang et al. (2023a) points out, we should be more cautious about using LLMs to evaluate texts in non-Latin languages. Additionally, existing research mainly focuses on more traditional NLG tasks such as translation, summarization, and dialogue. However, there are many new scenarios in reality with different requirements and evaluation criteria. Research on low-resource languages and new task scenarios will provide a more comprehensive understanding of LLMs’ evaluation capabilities.
Diverse forms of human-LLM collaborative NLG evaluation. According to the literature reviewed above, there is little research on collaborative evaluation between humans and LLMs. Neither humans nor LLMs are perfect, and each has its strengths. Since the ultimate goal of NLG research is to evaluate text quality more accurately and efficiently, we believe that collaboration between humans and LLMs can achieve better results than pure human evaluation or automatic evaluation. In the collaboration between humans and LLMs, technologies in the field of human-computer interaction may bring new implementation methods to the collaboration. In addition, what roles humans and LLMs should play in the evaluation and how they can better complement each other are still worth researching.